OpenStudy (anonymous):

solve by using distributive property 23 * 99 + 23

OpenStudy (cj49):

is it 23x99+23 or 23(99+23)

OpenStudy (anonymous):

Always multiply/divide first-then add/subtract, therefore its (23*99) + 23 =2300

OpenStudy (anonymous):

However, if the problem is written as 23(99+23), then using the distributive law would allow it to be (23*99) +(23*23) =2806

OpenStudy (anonymous):

where the number 23 is distributed to all numbers

jhonyy9 (jhonyy9):

so 23*99 +23 = 23(99+1) = 23*100 = 2300 hope so much that is understandably sure right easy